### v1

@@ 4 Global convergence
$d^{k}:=\mathbf{d}(x^{k})+\beta_{k}d^{k-1}+\frac{\beta_{k}\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k}),$ (4.3)
@@
Note that if $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k$, then it can be deduced form assumption $(H_{2})$ that the sequence $\{\Phi(x^{k},\mathbf{d}(x^{k}))\}$ is bounded. In fact, in this case, by the boundedness of $\{x^{k}\}$ and the continuity conclusions of Lemma 2.2, there exist $\gamma,\theta>0$ such that
$\|J\mathcal{F}(x^{k})\|\leq\gamma\quad\text{and}\quad\|\mathbf{d}(x^{k})\|\leq\theta$ (4.5)
for all $k$. Hence, using the continuity conclusions of Lemma 2.1, we can obtain the boundedness of $\{\Phi(x^{k},\mathbf{d}(x^{k}))\}$.
@@
Since $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, we have that $d^{k}\neq 0$ and $\Phi^{2}(x^{k},d^{k})/\|d^{k}\|^{2}$ is well defined. Since the step size $\alpha_{k}$ satisfies the generalized Wolfe conditions (LABEL:2*) for all $k\geq 0$, we can use [29, Proposition 3.3] to obtain the desired conclusion. ∎
@@
We present some auxiliary lemmas.
@@
By assumption $(H_{4})$, Lemma 2.2(ii) and Theorem 3.2, for any $k\geq 0$, we have
$\frac{\eta^{4}}{4\|d^{k}\|^{2}}\leq\frac{\|\mathbf{d}(x^{k})\|^{4}}{4\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},\mathbf{d}(x^{k}))}{\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},d^{k})}{\|d^{k}\|^{2}},$
which, using Lemma 4.1, yields that
@@
$d^{k}=\mathbf{d}(x^{k})+\|d^{k-1}\|\bigg{(}\beta_{k}u^{k-1}+\frac{\beta_{k}\Phi(x^{k},u^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
@@
$\|d^{k}\|=\mu_{k}(x^{k})+\nu_{k}(x^{k})\|d^{k-1}\|$ (4.6)
for all $k\geq 1$, where $\mu_{k},\nu_{k}:\mathbf{R}^{n}\to\mathbf{R}$. Using the fact $\|u^{k-1}\|=1$, the boundedness of $\{x^{k}\}$, and the continuity conclusions of Lemmas 2.1 and 2.3, we can infer that the sequences $\{\mu_{k}(x^{k})\}$ and $\{\nu_{k}(x^{k})\}$ are bounded. Hence, there exists $C_{\mu}>0$ such that
$|\mu_{k}(x^{k})|\leq C_{\mu}$
for all $k\geq 1$. By the boundedness of $\{\nu_{k}(x^{k})\}$, we obtain that $\liminf_{k\to\infty}\nu_{k}(x^{k})$ exists. Moreover, we can assert that
@@
Otherwise, it follows from (4.6) that
$\liminf_{k\to\infty}\|d^{k}\|<\liminf_{k\to\infty}\mu_{k}(x^{k})\leq C_{\mu}<\infty,$
which contrary to Lemma 4.2. Based on this, we divide the sequence $\{k\}$ into the two subsequences
@@
By the definition of the limit inferior, the sequence $\{\underline{k}\}$ is finite. According to Lemma 4.2, there exists $N_{0}>0$ such that for any $k\geq N_{0}$,
$\|d^{k}\|>\frac{2C_{\mu}+1}{\omega}.$ (4.7)
Let
$\varpi^{k}:=\frac{1}{\|d^{k}\|}\bigg{(}\mathbf{d}(x^{k})+\frac{\beta_{k}\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
@@
By (3.3), (4.6), (4.7), the continuity conclusions of Lemmas 2.1 and 2.2, the boundedness of $\{x^{k}\}$, and Lemmas 4.1 and 4.2, we obtain from (4.9) the following estimate
@@
$\leq\sum^{\infty}_{k=1}\frac{4}{\|d^{k}\|^{2}}\bigg{(}\|\mathbf{d}(x^{k})\|+\frac{\beta_{k}|\Phi(x^{k},d^{k-1})|}{|\Phi(x^{k},-\mathbf{d}(x^{k}))|}\|\mathbf{d}(x^{k})\|\bigg{)}^{2}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum^{\infty}_{k=1}\frac{8\beta^{2}_{k}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{k})\|^{2}}{\Phi^{2}(x^{k},-\mathbf{d}(x^{k}))}\frac{\Phi^{2}(x^{k-1},d^{k-1})}{\|d^{k}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},-\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{32\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},-\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\omega^{2}\|d^{\overline{k}-1}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},-\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
@@
Here, we use the finiteness of the sequences $\{\overline{k}\leq N\}$ and $\{\underline{k}\}$, and the proof is complete.
@@
Suppose, contrary to our claim, that assumption $(H_{4})$ holds. By assumption $(H_{3})$, Lemmas 2.1(iii) and 2.2(ii), and (4.5), it follows from (4.6) that
@@
$\leq\frac{2L\|\mathbf{d}(x^{k})\|\|x^{k}-x^{k-1}\|}{|\mathbf{d}(x^{k-1})|^{2}}$
@@
Moreover, it can be inferred from Lemma 2.2(ii) and (4.5) that
@@
where $C_{1}:\max\{1,\sigma/(1-\sigma)\}$. From (3.6) and Lemma 2.2(ii), we have
$|\Phi(x^{k},-\mathbf{d}(x^{k}))|\geq|\Phi(x^{k},\mathbf{d}(x^{k}))|>\frac{\|\mathbf{d}(x^{k})\|^{2}}{2}\geq\frac{\eta^{2}}{2}.$ (4.13)
Combining (4.5), (4.10), (4.11), (4.12) and (4.13), we infer from (4.3) that
$\|d^{k}\|^{2}$ $\leq\bigg{(}\|\mathbf{d}(x^{k})\|+\beta_{k}\|d^{k-1}\|+\frac{\beta_{k}|\Phi(x^{k},d^{k-1})|\|\mathbf{d}(x^{k})\|}{|\Phi(x^{k},-\mathbf{d}(x^{k}))|}\bigg{)}^{2}$ (4.14)
@@
By the definition of $u^{k}$, we deduce that for any $m>m_{0}$,
@@
$=\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|u^{m_{0}}+\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|(u^{i-1}-u^{m_{0}}),$
where $s^{i-1}:=x^{i}-x^{i-1}$. From this, we get
@@
for all $N\geq\lceil 8CM\rceil$. Combining (4.18) and (4.19), for any $m\geq m_{0}+\lceil 8CM\rceil$, we have
$\|d^{m}\|^{2}$ $\leq 2{\theta}^{2}\bigg{(}\sum^{m_{0}+\lceil 8CM\rceil}_{i=m_{0}+1}\prod^{m_{0}+\lceil 8CM\rceil-1}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}$ (4.20)
@@ 5 Numerical experiments
The considered methods in the numerical tests are as follows:
  • TT-PRP CG method: The conjugate parameter and the step size are given by Algorithm TT-PRP.
  • TT-PRP1 CG method: The conjugate parameter is defined by (4.3), and the step size satisfies the strong Wolfe conditions (2.6).
  • PRP+ CG method [29]: The conjugate parameter is given by (3.2), and the step size is obtained by a strong Wolfe line search procedure.
  • The steepest descent (SD) method [15]: The step size is governed by the strong Wolfe conditions (2.6).
In our implementation, the parameters in Definitions 2.3 and 2.4 are chosen as $\rho=0.01$, $\sigma=0.2$ and $\lambda=2$, and the maximum number of allowed iterations is set to 3000.
@@
evalf 53.0 55.0 55.0 465.5 evalf 34.0 37.0 37.0 60.0
evalg 43.0 46.0 46.0 398.0 evalg 26.0 29.0 29.0 50.0
@@
it 33.0 32.5 38.0 70.5 it 134.0 145.0 183.5 214.0
evalf 276.0 308.0 375.5 523.5 evalf 738.0 964.0 1287.0 1475.0
evalg 242.0 273.0 337.5 451.0 evalg 602.0 837.0 1101.0 1258.0
@@
it 6.0 6.0 6.0 8.0 it 151.0 163.0 220.0 283.0
evalf 52.0 61.0 61.0 82.0 evalf 784.0 960.5 1682.5 2264.5
evalg 44.0 53.0 53.0 72.0 evalg 611.0 795.0 1460.5 1979.5
@@
it 130.0 145.5 171.5 221.5 it 3.0 3.0 3.0 3.0
evalf 779.5 1163.5 1363.5 2687.5 evalf 28.0 28.0 28.0 28.0
evalg 911.5 1311.0 1537.0 2464.0 evalg 23.0 23.0 23.0 23.0
@@
it 8.0 9.0 10.0 16.5 it 8.0 8.0 8.0 27.5
evalf 55.0 72.0 78.0 102.5 evalf 47.0 58.0 58.0 165.0
evalg 45.0 61.0 66.0 84.0 evalg 37.0 48.0 48.0 135.0
@@
it 3.0 3.0 3.0 3.0 it 20.5 22.0 32.5 62.0
evalf 26.0 26.0 26.0 26.0 evalf 301.0 328.5 517.0 928.5
evalg 21.0 21.0 21.0 21.0 evalg 278.5 304.5 482.5 864.5
@@
it 1.0 1.0 1.0 1.0 it 20.0 21.5 33.0 58.0
evalf 6.0 6.0 6.0 6.0 evalf 212.0 241.0 360.5 772.0
evalg 5.0 5.0 5.0 5.0 evalg 189.0 217.5 325.5 712.0
The ability of various algorithms to successfully obtain Pareto frontiers has always attracted great interest in vector optimization problems with finite objectives. To evaluate the effectiveness of the TT-PRP method in generating Pareto frontiers, the value spaces of the bicriteria problems (AP3, Far1 and Hil1) and the three-criteria problems (FDS-1, MOP5 and MOP7) are plotted in Figures 1 and 2 using the TT-PRP method. For these problems, 400 starting points are randomly generated (except for problem Far1, which uses 600 points due to the complexity of the image space) in the box given in the column “$x_{D}$” in Table 1. In the figures, a green point represents a final iteration, while the beginning of a straight segment represents the corresponding starting point. The results in Figures 1 and 2 show the effective estimation of the Pareto frontier for the considered problems using a reasonable number of initial points by the TT-PRP method.

### v2

@@ 4 Global convergence
$d^{k}:=\mathbf{d}(x^{k})+\beta_{k}d^{k-1}-\beta_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k}),$ (4.3)
@@
Note that if $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, then it can be deduced form assumption $(H_{2})$ that the sequence $\{\Phi(x^{k},\mathbf{d}(x^{k}))\}$ is bounded. In fact, in this case, by the boundedness of $\{x^{k}\}$ and the continuity conclusions of Lemma 2.3, there exist $\gamma,\theta>0$ such that
$\|J\mathcal{F}(x^{k})\|\leq\gamma\quad\text{and}\quad\|\mathbf{d}(x^{k})\|\leq\theta,$ (4.5)
which, along with the definitions of $\Phi$ and $V$, yields that
$\|\Phi(x^{k},\mathbf{d}(x^{k}))\|=|\sup\{\langle\mathcal{J}F(x^{k})\mathbf{d}(x^{k}),v\rangle:v\in V\}|$ (4.6)
$\leq\sup\{\|J\mathcal{F}(x^{k})\mathbf{d}(x^{k})\|\|v\|:v\in V\}\leq\|J\mathcal{F}(x^{k})\|\|\mathbf{d}(x^{k})\|\leq\gamma\theta.$
We next show that the iterative sequence generated by Algorithm TT-PRP satisfies the Zoutendijk condition.
@@
Since $d^{k}$ is an $E$-descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, we have that $d^{k}\neq 0$ and $\Phi^{2}(x^{k},d^{k})/\|d^{k}\|^{2}$ is well defined. Since the step size $\alpha_{k}$ satisfies the generalized Wolfe conditions (LABEL:2*) for all $k\geq 0$, we can use [36, Proposition 3.3] to obtain the desired conclusion. ∎
@@
Based on the contradictory assumption, we have the following auxiliary lemmas.
@@
By assumption $(H_{4})$, Lemma 2.3(ii) and Theorem 3.2, we have
$\frac{\eta^{4}}{4\|d^{k}\|^{2}}\leq\frac{\|\mathbf{d}(x^{k})\|^{4}}{4\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},\mathbf{d}(x^{k}))}{\|d^{k}\|^{2}}\leq\frac{\Phi^{2}(x^{k},d^{k})}{\|d^{k}\|^{2}}$
for all $k\geq 0$, which, using Lemma 4.1, yields that
@@
$d^{k}=\|d^{k-1}\|\bigg{(}\frac{\mathbf{d}(x^{k})}{\|d^{k-1}\|}+\beta_{k}u^{k-1}-\beta_{k}\frac{|\Phi(x^{k},u^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
@@
$\|d^{k}\|=\nu_{k}(x^{k})\|d^{k-1}\|$ (4.7)
for all $k\geq 1$, where
$\nu_{k}:=\bigg{\|}\frac{\mathbf{d}(x^{k})}{\|d^{k-1}\|}+\beta_{k}u^{k-1}-\beta_{k}\frac{|\Phi(x^{k},u^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{\|}.$
Using the fact $\|u^{k-1}\|=1$, the boundedness of $\{x^{k}\}$, the boundedness of $\{\frac{1}{\|d^{k-1}\|}\}$ obtained from Lemma 4.2 and the continuity conclusions of Lemmas 2.2 and 2.3, we can infer that the sequence $\{\nu_{k}(x^{k})\}$ is bounded, then $\liminf_{k\to\infty}\nu_{k}(x^{k})$ exists. Moreover, we can assert from (4.7) that
@@
Based on this, we divide the sequence $\{k\}$ into the two subsequences
@@
By the definition of the limit inferior, the sequence $\{\underline{k}\}$ is finite. Let
$\varpi^{k}:=\frac{1}{\|d^{k}\|}\bigg{(}\mathbf{d}(x^{k})-\beta_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$
@@
From Lemma 2.3(ii) and assumption $(H_{4})$, we have
$|\Phi(x^{k},\mathbf{d}(x^{k}))|>\frac{\|\mathbf{d}(x^{k})\|^{2}}{2}\geq\frac{\eta^{2}}{2}.$ (4.10)
By (4.1b), (4.7), (4.10), the continuity conclusions of Lemmas 2.2 and 2.3, the boundedness of $\{x^{k}\}$, and Lemmas 4.1 and 4.2, we obtain from (4.9) the following estimate
@@
$\leq\sum^{\infty}_{k=1}\frac{4}{\|d^{k}\|^{2}}\bigg{\|}\mathbf{d}(x^{k})-\beta_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{\|}^{2}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum^{\infty}_{k=1}\frac{8\beta^{2}_{k}\Phi^{2}(x^{k},d^{k-1})\|\mathbf{d}(x^{k})\|^{2}}{\Phi^{2}(x^{k},\mathbf{d}(x^{k}))\|d^{k}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum^{\infty}_{k=1}\frac{8\beta^{2}_{k}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{k})\|^{2}}{\Phi^{2}(x^{k},\mathbf{d}(x^{k}))}\frac{\Phi^{2}(x^{k-1},d^{k-1})}{\|d^{k}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
$\leq\sum^{\infty}_{k=1}\frac{8\|\mathbf{d}(x^{k})\|^{2}}{\|d^{k}\|^{2}}+\sum_{\overline{k}\leq N_{0}}\frac{8\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\Phi^{2}(x^{\overline{k}},\mathbf{d}(x^{\overline{k}}))}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\|d^{\overline{k}}\|^{2}}$
$\quad+\sum_{\overline{k}>N_{0}}\frac{32\beta^{2}_{\overline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\overline{k}})\|^{2}}{\eta^{4}}\frac{\Phi^{2}(x^{\overline{k}-1},d^{\overline{k}-1})}{\omega^{2}\|d^{\overline{k}-1}\|^{2}}$
$\quad+\sum_{\underline{k}}\frac{8\beta^{2}_{\underline{k}}\max\{\lambda^{2},\sigma^{2}\}\|\mathbf{d}(x^{\underline{k}})\|^{2}}{\Phi^{2}(x^{\underline{k}},\mathbf{d}(x^{\underline{k}}))}\frac{\Phi^{2}(x^{\underline{k}-1},d^{\underline{k}-1})}{\|d^{\underline{k}}\|^{2}}$
@@
Here, we use the finiteness of the sequences $\{\overline{k}\leq N\}$ and $\{\underline{k}\}$, and the proof is complete. ∎
We now proceed to establish the global convergence of Algorithm TT-PRP.
@@
Suppose, contrary to our claim, that assumption $(H_{4})$ holds. By assumption $(H_{3})$, Lemmas 2.2(iii) and 2.3(ii), and (4.5), it follows from (4.7) that
@@
$\leq\frac{2L\|\mathbf{d}(x^{k})\|\|x^{k}-x^{k-1}\|}{\|\mathbf{d}(x^{k-1})\|^{2}}$
@@
Moreover, it can be inferred from Lemma 2.3(ii), (4.6) and (4.10) that
@@
where $C_{1}:\max\{1,\sigma/(1-\sigma)\}$. Combining (4.5) with (4.10)-(4.13), we infer from (4.3) that
$\|d^{k}\|^{2}$ $\leq\bigg{(}\|\mathbf{d}(x^{k})\|+\beta_{k}\|d^{k-1}\|+\frac{\beta_{k}|\Phi(x^{k},d^{k-1})|\|\mathbf{d}(x^{k})\|}{|\Phi(x^{k},\mathbf{d}(x^{k}))|}\bigg{)}^{2}$ (4.14)
@@
$\leq\bigg{(}\theta+\frac{2L\eta^{2}+8C_{1}L\gamma\theta^{2}}{\eta^{4}}\|x^{k}-x^{k-1}\|\|d^{k-1}\|\bigg{)}^{2}$
@@
By the definition of $u^{k}$, we deduce that
@@
$=\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|u^{m_{0}}+\sum^{m}_{i=m_{0}+1}\|s^{i-1}\|(u^{i-1}-u^{m_{0}})$
for all $m>m_{0}$, where $s^{i-1}:=x^{i}-x^{i-1}$. From this, we get
@@
for all $N\geq\lceil 8CM\rceil$. Combining (4.18) with (4.19), for any $m\geq m_{0}+\lceil 8CM\rceil$, we have
$\|d^{m}\|^{2}$ $\leq 2{\theta}^{2}\bigg{(}\sum^{m_{0}+\lceil 8CM\rceil}_{i=m_{0}+1}\prod^{m_{0}+\lceil 8CM\rceil-1}_{j=i}2C^{2}\|s^{j-1}\|^{2}\bigg{)}$
@@ 5 Numerical experiments
We now show the effectiveness and characteristics of the proposed three-term CG algorithm in comparison with traditional algorithms. The considered methods in the numerical tests are as follows:
  • TT-PRP: The conjugate parameter and the step size are given by Algorithm TT-PRP.
  • TT-PRP1: The conjugate parameter is defined by (4.3), and the step size satisfies the strong Wolfe conditions (2.6).
  • PRP+ [36]: The conjugate parameter is given by (3.2), and the step size is obtained by a strong Wolfe line search procedure.
  • Steepest descent (SD) [19]: The step size is governed by the strong Wolfe conditions (2.6).
In our implementation, the parameters in Definitions 2.5 and 2.6 are chosen as $\rho=0.01$, $\sigma=0.2$ and $\lambda=2$, and the maximum number of allowed iterations is set to 3000.
@@
evalf 46.0 51.0 55.0 465.5 evalf 26.0 31.0 37.0 60.0
evalg 37.0 42.0 46.0 398.0 evalg 18.0 23.0 29.0 50.0
@@
it 33.0 32.5 38.0 70.5 it 93.0 98.0 183.5 214.0
evalf 276.0 308.0 375.5 523.5 evalf 517.5 642.0 1287.0 1475.0
evalg 242.0 273.0 337.5 451.0 evalg 422.5 542.0 1101.0 1258.0
@@
it 6.0 6.0 6.0 8.0 it 134.5 141.5 220.0 283.0
evalf 51.0 59.0 61.0 82.0 evalf 713.5 840.5 1682.5 2264.5
evalg 43.0 51.0 53.0 72.0 evalg 577.0 697.0 1460.5 1979.5
@@
it 122.5 139.0 171.5 221.5 it 2.0 2.0 3.0 3.0
evalf 735.0 1266.5 1537.0 2687.5 evalf 19.0 21.0 28.0 28.0
evalg 859.5 1311.0 1125.5 2464.0 evalg 15.0 17.0 23.0 23.0
@@
it 6.5 7.0 10.0 16.5 it 7.0 7.0 8.0 27.5
evalf 38.0 56.0 78.0 102.5 evalf 36.5 48.0 58.0 165.0
evalg 29.5 47.0 66.0 84.0 evalg 27.5 39.0 48.0 135.0
@@
it 2.0 2.0 3.0 3.0 it 18.0 18.5 32.5 62.0
evalf 18.0 18.0 26.0 26.0 evalf 238.0 303.5 517.0 928.5
evalg 14.0 14.0 21.0 21.0 evalg 218.0 283.0 482.5 864.5
@@
it 1.0 1.0 1.0 1.0 it 18.5 20.0 33.0 58.0
evalf 6.0 6.0 6.0 6.0 evalf 230.0 336.0 360.5 772.0
evalg 5.0 5.0 5.0 5.0 evalg 209.5 314.0 325.5 712.0
The ability of various algorithms to successfully approximate Pareto frontiers (Pareto optimal or weak Pareto optimal solutions) has always attracted great interest in vector optimization problems with finite objectives. To evaluate the effectiveness of the TT-PRP method in generating Pareto frontiers, the value spaces of the bicriteria problems (AP3, Far1 and Hil1) and the three-criteria problems (FDS-1, MOP5 and MOP7) are plotted in Figures 1 and 2 using the TT-PRP method. For these problems, 400 starting points are randomly generated (except for problem Far1, which uses 600 points due to the complexity of the image space) in the box given in the column “$x_{D}$” in Table 1. In the figures, a green point represents a final iteration, while the beginning of a straight segment represents the corresponding starting point. The results in Figures 1 and 2 show the effective estimation of the Pareto frontier for the considered problems using a reasonable number of initial points by the TT-PRP method.